The manager of Perrotti’s Pizza collects data concerning customer complaints. Using the information below, draw a Pareto chart to identify the critical few problems. Based on the completed Pareto chart, what is the cumulative percentage for the first two most critical issues? Problem Frequency Topping is stuck to box lid 17 Pizza arrives late 35 Wrong topping or combinations 9 Wrong style of crust 6 Wrong size 4 Pizza is partially eaten 2 Pizza never arrives 8 Q14. The cumulative percentage for the single most critical issue is ______ percent. (Enter your response as a nearest integer.) Q15. The cumulative percentage for the first two most critical issues is ______ percent. (Enter your response as a nearest
Answer #1

Sum of frequency = 17+35+9+6+4+2+8 = 81

Sum of frequency of the two most critical issues = 35+17 = 52

Answer 14: Cumulative percentage for the single most critical issues = 35/81= 43.2 % ~43%

Answer 15: Cumulative percentage for the first two most critical issues = 52/81= 64.2 % ~64%